# A Collaborative Content Moderation Framework for Toxicity Detection based on Conformalized Estimates of Annotation Disagreement

### v2

@@ Abstract
Content moderation typically combines the efforts of human moderators and machine learning models. However, these systems often rely on data where significant disagreement occurs during moderation, reflecting the subjective nature of toxicity perception. Rather than dismissing this disagreement as noise, we interpret it as a valuable signal that highlights the inherent ambiguity of the content—an insight missed when only the majority label is considered. In this work, we introduce a novel content moderation framework that emphasizes the importance of capturing annotation disagreement. Our approach uses multitask learning, where toxicity classification serves as the primary task and annotation disagreement is addressed as an auxiliary task. Additionally, we leverage uncertainty estimation techniques, specifically Conformal Prediction, to account for both the ambiguity in comment annotations and the model’s inherent uncertainty in predicting toxicity and disagreement. The framework also allows moderators to adjust thresholds for annotation disagreement, offering flexibility in determining when ambiguity should trigger a review. We demonstrate that our joint approach enhances model performance, calibration, and uncertainty estimation, while offering greater parameter efficiency and improving the review process in comparison to single-task methods.
@@
Conformal prediction , Uncertainty quantification , Collaborative content moderation , Toxicity detection
journal: Pattern Recognition
@@
We propose using a multitask approach with auxiliary task annotation disagreement to guide reviews.
@@ 1 Introduction
Although incorporating uncertainty has advanced ML models in different disciplines, such as traffic modelling [6] or medical imaging [7], it is crucial to recognize that tasks like toxic language detection, which are inherently subjective, require a different perspective due to their complexity. Disagreements among annotators during labelling processes often occur in content moderation reflecting this subjectivity, which can lead to bias and, in some cases, to the censorship of minority opinions. The quantification of inter-annotator disagreement is thus a valuable source of information about the complexity inherent in comments, and is directly related to the aleatoric uncertainty present in the large databases whose inputs require moderation.
Annotation disagreement quantification, together with the epistemic uncertainty of the modelling process, allow us to identify the cases when the model cannot provide a robust prediction, as well as those when, due to their complexity and subjectivity, the intervention of a human moderator is necessary. Accounting for both cases is critical, as ML models trained with a majority vote disregard the inherent ambiguity of comments. Unfortunately, considering different annotation perspectives alongside the data modelling pipeline has been observed not to scale well in diverse annotation processes [8]. Additionally, as toxic language evolves over time, it is important to endow content moderators with the tools to make final decisions.
This manuscript builds upon the above observations. Specifically, we propose a novel modelling framework that unifies the primary task of toxic language classification with the auxiliary task of annotation disagreement quantification through a multitask framework architecture. The proposed approach is augmented with UQ for both tasks, using Inductive Conformal Prediction (CP) for this purpose. This proposal is grounded on a previous study by Fornaciari et al. [9], which observed that unifying annotation disagreement and classification can be beneficial for the performance of the primary task. Finally, it aims to address a need highlighted by Gillespie [10], who stated: “Perhaps, automated tools are best used to identify the bulk of the cases, leaving the less obvious or more controversial identifications to human reviewers, as societies hold together not by reaching perfect consensus, but by keeping their values under constant and legitimate reconsideration”.
@@
These three research questions aim to analyze the improvements the multitask approach yields compared to simple single-task models, or, in the case of the full framework, compared to performing quantification separately using a composite of single-task models (CoM).
@@ 2 Related Work and Contribution
In this section, we first establish the fundamentals and perform a literature review of CP (Subsection 2.1), with a focus on text classification. Then we define content moderation and revisit frameworks proposed in the recent literature within the crossroads of toxicity detection and uncertainty estimation (Subsection 2.2). Finally, we end the section with a statement on the contribution of our work beyond the literature reviewed in the previous subsections (Subsection 2.3).
@@ 2.1 Conformal Prediction
Given its robust theoretical foundation and model-agnostic applicability, CP has been widely utilized in various domains, including object detection, medicine, and natural language processing (NLP). According to a recent survey on the use of CP in NLP [13], tasks often tackled with CP include binary classification, particularly for paraphrase detection, sentiment analysis, and Boolean question answering [14]; multiclass classification for part-of-speech (POS) tagging [15]; and multilabel classification [16]. In the case of natural language generation, CP has been used at the sentence level to provide guarantees for Large Language Model (LLM) planners in completing tasks [17], and at the token level to mitigate overconfidence in nucleus sampling text decoding strategies [18]. Although CP has been predominantly applied to classification tasks, we have not observed any instances in the literature where it has been specifically used for toxic language detection, except for the introductory examples provided in Angelopoulos and Bates [19]. This lack of research is evident in the next subsection, where we explore content moderation in further detail.
@@ 2.2 Content Moderation
Before the rise of social media, CM was primarily performed by human moderators. However, with the surge of interactions on new platforms, this approach became unfeasible for humans to manage alone. As a result, companies like YouTube, Twitter, and Facebook began to develop their own moderation tools. Other initiatives also emerged, such as Google’s Perspective API and OpenAI’s CM tools, to support companies that did not have the resources to create their own tools. These algorithmic moderation tools were primarily designed to automate CM, often without fully considering the role of human moderators. Moderators were left mainly to review content that was unjustifiably censored due to biases in these models. This was shown by Hutchinson et al. [21] where instances of biases in Perspective API are very clearly outlined.
@@ 2.3 Main Contribution
Taking into account the analysed literature, the main contribution of this manuscript is a collaborative CM framework where the moderator is in control of the algorithmic CM tools by deciding when the moderator’s final decision should be incorporated. This is achieved not only by considering the model’s confidence through UQ, but also by incorporating information about the level of annotation disagreement the model should handle in its predictions. Since toxicity is a subjective task, with opinions varying across content annotators, it is crucial to leave the final judgment to those who are familiar with CM policies in place, as they are the ones who best understand these policies and their context of application. Additionally, given the evolving nature of language, it is of utmost importance to have mechanisms that are aware of this variability, especially in what refers to the ambiguity of textual toxicity. Along this line, the collaborative CM framework proposed in this work introduces information about the annotation disagreement into the modelling process within a multitask learning setting, enriching the conformalized estimation of the model’s confidence, and ultimately yielding better and more reliable decisions about the toxicity of the text data fed at its input. As our experiments in Section 5 will clearly show, our multitask architecture yields competitive levels of detection performance, a more aligned and better calibrated estimation of the model’s confidence, and expanded information (predicted annotation ambiguity) that allows for more informed decisions about the moderation of potentially toxic content.
@@ 3.1 Dataset
The Jigsaw Unintended Bias in Toxicity Classification dataset, created as part of a challenge organized by Google Jigsaw, provides a high-quality dataset that is already divided into training, validation, and test sets. This predefined partitioning facilitates the reproducibility of our experiments. Additionally, compared to other datasets in the toxicity domain [4], this dataset features a substantially larger group of annotators (8,899), as opposed to the smaller groups of tens typically found in other datasets. This large number of annotators is a significant advantage, as it provides a more diverse range of opinions, not only in terms of numbers, but also in terms of background, ethnicity, gender, and age. However, during the dataset preprocessing phase, the following caveats are found:
@@ 3.2 Methods
In this section we introduce the notation, selected model architecture, learning objectives, and hyperparameters (Subsection 3.2.1), whereas Subsection 3.2.2 discusses various methods from the literature for computing disagreement and explain our selected approach. Finally, Subsection 3.2.3 presents the CP algorithms considered in our experiments.
3.2.1 Notation, Model Selection, Learning Objectives and Hyperparameters
@@
As depicted in Figure 1, the MTL CM process consists of a primary task (toxicity classification of the given comment $\mathbf{x}_{i}$) and an auxiliary task (estimation of the annotation disagreement). The architecture is based on $\text{DistillBERT}_{\text{base uncased}}$ [27], which serves to extract the embedding from the CLS token (a sentence-level representation for classification) of the given comment $\mathbf{x}_{i}$ to perform the MTL task. The generated predictions for each task are defined as $\widehat{y}_{i}$ from the primary task and $\widehat{d}_{i}(\mathbf{x}_{i})$ from the auxiliary task. We then generate the corresponding prediction set $\mathcal{C}(\mathbf{x}_{i})$ and confidence interval $I(\mathbf{x}_{i})$ using the inductive CP techniques later detailed in Subsection 2.1 and a calibration subset $\mathcal{D}_{\text{cal}}$. Finally, the need for a human moderator’s revision is determined by whether prediction set length of $\mathcal{C}(\mathbf{x}_{i})$ is 2 (namely, the number of possible classes in our problem) or if the predicted confidence interval $I(\mathbf{x}_{i})$ is larger than an ambiguity threshold $\gamma$. This criterion allows considering the model’s inherent uncertainty in the predictions of the primary and auxiliary tasks, as well as the ambiguity of the comments.
@@ 3.2.3 Conformal Algorithms under Consideration
  • Class-Conditional LAC (CCLAC) [19]: In this case we compute the conformity score $s(\mathbf{x},y)$ for each sample $\mathbf{x}$ and class $y$ as one minus the softmax probability assigned by the model to the class, either toxic or non-toxic. The quantile is then computed for each class, yielding a quantile value $\widehat{q}_{y}$ for each class $y$ (toxic and not toxic). Finally, during inference we iterate through the classes and include them in the prediction set associated to query instance $\mathbf{x}_{i}$ if their conformity score $s(\mathbf{x}_{i},y)$ is below the corresponding quantile $\widehat{q}_{y}$, namely:
@@
    $I(\mathbf{x}_{i})=\left[\widehat{d}_{i}(\mathbf{x}_{i})\cdot(1-\widehat{q}),\widehat{d}_{i}(\mathbf{x}_{i})\cdot(1+\widehat{q})\right].$ (7)
@@
    $I(\mathbf{x}_{i})=\left[d\left(\min\left\{b\in\{1,\ldots,B\}:\;s(\mathbf{x}_{i},d(b))\geq\widehat{q}\right\}\right)\right.,$
    $\left.d\left(\max\left\{b\in\{1,\ldots,B\}:\;s(\mathbf{x}_{i},d(b))\geq\widehat{q}\right\}\right)\right].$ (10)

### v3

# A Collaborative Content Moderation Framework for Toxicity Detection based on Multitask Neural Networks and Conformal Estimates of Annotation Disagreement

@@ Abstract
Content moderation typically combines the efforts of human moderators and machine learning models. However, these systems often rely on data where significant disagreement occurs during moderation, reflecting the subjective nature of toxicity perception. Rather than dismissing this disagreement as noise, we interpret it as a valuable signal that highlights the inherent ambiguity of the content—an insight missed when only the majority label is considered. In this work, we introduce a novel content moderation framework that emphasizes the importance of capturing annotation disagreement. In this work, we propose a novel content moderation framework that prioritizes capturing annotation disagreement. Our approach leverages multitask neural networks with transformer architectures as their backbone, where toxicity classification serves as the primary task and annotation disagreement is modelled as an auxiliary task. By framing disagreement as a predictive problem within the multitask learning architecture, our method effectively captures the nuanced ambiguity of content toxicity. Additionally, we leverage uncertainty estimation techniques, specifically Conformal Prediction, to account for the model’s inherent uncertainty in predicting toxicity and annotation disagreement. The framework also allows moderators to adjust thresholds for annotation disagreement, offering flexibility in determining when ambiguity should trigger a review. We demonstrate that our joint approach enhances model performance, calibration, and uncertainty estimation, while offering greater parameter efficiency and improving the review process in comparison to single-task methods.
@@
Conformal prediction , Uncertainty quantification , Multitask neural network , Collaborative content moderation , Toxicity detection
journal: Neurocomputing
@@
A multitask neural network with auxiliary annotation disagreement task to guide reviews.
@@ 1 Introduction
Incorporating uncertainty has advanced ML models in different disciplines, such as traffic modelling [6] or medical imaging [7]. This advance has been particularly notable in deep neural networks, in which UQ is nowadays crucial for the trustworthiness of such opaque learning models [8]. However, it is crucial to recognize that tasks like toxic language detection, which are inherently subjective, require a different perspective due to their complexity. Disagreements among annotators during labelling processes often occur in content moderation reflecting this subjectivity, which can lead to bias and, in some cases, to the censorship of minority opinions. The quantification of inter-annotator disagreement is thus a valuable source of information about the complexity inherent in comments, and is directly related to the aleatoric uncertainty present in the large databases whose inputs require moderation.
Quantifying annotation disagreement and epistemic uncertainty in the modelling process allows us to identify cases where the model cannot provide robust predictions and must refrain from making them [9]. Additionally, it helps determine situations where human moderation is necessary due to complexity or subjectivity. Accounting for both cases is critical, as ML models trained with a majority vote disregard the inherent ambiguity of comments. Unfortunately, considering different annotation perspectives alongside the data modelling pipeline has been observed not to scale well in diverse annotation processes [10]. Additionally, as toxic language evolves over time, it is important to endow content moderators with the tools to make final decisions.
This manuscript builds upon the above observations. Specifically, we propose a novel modeling framework that unifies the primary task of toxic language classification with the auxiliary task of annotation disagreement quantification through a multitask learning architecture approach leveraging a BERT-based Transformer network. The proposed approach is augmented with UQ for both tasks, using Inductive Conformal Prediction (CP) for this purpose. This proposal is grounded on a previous study by Fornaciari et al. [11], which observed that unifying annotation disagreement and classification can be beneficial for the performance of the primary task. Finally, it aims to address a need highlighted by Gillespie [12], who stated: “Perhaps, automated tools are best used to identify the bulk of the cases, leaving the less obvious or more controversial identifications to human reviewers, as societies hold together not by reaching perfect consensus, but by keeping their values under constant and legitimate reconsideration”.
@@
These three research questions aim to analyse the improvements the multitask approach yields compared to simple single-task models, or, in the case of the full framework, compared to performing quantification separately using a composite of single-task models (CoM).
@@ 2 Related Work and Contribution
In this section, we first establish the fundamentals and perform a literature review of CP (Subsection 2.1), with a focus on text classification. Furthermore, we review the research area of perspectivism in Section 2.2, with a particular focus on learning from annotation disagreement. Then we define content moderation and revisit frameworks proposed in the recent literature within the crossroads of toxicity detection and uncertainty estimation (Subsection 2.3). Finally, we end the section with a statement on the contribution of our work beyond the literature reviewed in the previous subsections (Subsection 2.4).
@@ 2.1 Conformal Prediction
Given its robust theoretical foundation and model-agnostic applicability, CP has been widely utilized in various domains, including object detection [15], medicine [16], weather forecast [17] and natural language processing (NLP). According to a recent survey on the use of CP in NLP [18], tasks often tackled with CP include binary classification, particularly for paraphrase detection, sentiment analysis, and Boolean question answering [19]; multiclass classification for part-of-speech (POS) tagging [20]; and multilabel classification [21]. In the case of natural language generation, CP has been used at the sentence level to provide guarantees for Large Language Model (LLM) planners to complete tasks [22], and at the token level to mitigate overconfidence in nucleus sampling text decoding strategies [23]. Although CP has been predominantly applied to classification tasks, we have not observed any instances in the literature where it has been specifically used for toxic language detection, except for the introductory examples provided in Angelopoulos and Bates [24]. This lack of research is evident in the next subsection, where we explore content moderation in more detail.
2.2 Perspectivism: Learning from annotation disagreement
In supervised learning it is common practice to collect multiple annotator judgments on the same data instances to improve the quality of the final labels [25]. Although all human judgments carry some degree of subjectivity, certain tasks, including toxicity detection, are widely acknowledged to be particularly subjective [26]. A commonly used approach to fuse different annotations for a given data instance is majority voting, which is widely adopted in the literature when such annotations are available. However, this fusion approach may not always be appropriate. Such aggregation methods often overlook the socially constructed nature of human perceptions, silencing valuable insights especially those from underrepresented groups [27]. This is particularly problematic because these methods fail to fully capture the diversity of opinions, leading to the loss of important information especially in tasks where a unique ’ground truth’ may not even exist [28].
Although annotator disagreement has been identified as a meaningful indicator in these subjective tasks, it is equally important to recognize that such disagreement does not always reflect genuine differences in perception. In many cases, it can arise due to issues in the annotation process, such as sloppy annotations, ambiguity, or missing information, which can negatively impact the quality of these disagreement signals [29]. These factors have been widely documented in studies that analyse disagreement as a positive cue, and are commonly considered to be an inherent limitation of these approaches [30, 31, 32] due to the difficulty in detecting them [33]. Given this limitation, Cabitza et al. [34] and Aroyo and Welty [33] have proposed to introduce improved annotation procedures. Key factors highlighted in these works include ensuring a sufficient number of annotators (at least 12 to 15 per data instance) to allow for statistically significant majorities. This recommendation is based on the finding that beyond this number, significant differences in disagreement are rarely observed. In addition, they emphasize the importance of involving a diverse set of raters in terms of origin, culture, background, and other relevant aspects to better capture the complexity of subjective tasks.
The challenge of effectively handling disagreement in annotation leads to the emergence of perspectivism, a theoretical framework and a family of methods in Artificial Intelligence (AI) that takes into account the importance of annotation disagreement for predictive models and embraces the use of labelled data where disaggregated labels are available [35]. There are two categories within perspectivism: weak perspectivism, which focuses on the collection and analysis of disaggregated labels; and strong perspectivism, where disaggregated labels are used for both training and evaluation. In this work we focus on strong perspectivism, particularly in learning from disagreement approaches, where we primarily observe three main methodologies: i) modelling individual perspectives, ii) modelling disagreement among groups of annotators, and iii) modelling label distribution.
To begin with, modelling individual perspectives focuses on learning from each annotator’s labels in isolation, rather than aggregating all annotations under a majority voting rule. All existing approaches in the literature share a common procedure: incorporating an annotator embedding in addition to the text embedding to personalize the model’s predictions. Yin et al. [36] first embraced this methodology. Other approaches, besides including annotator embeddings, also incorporate annotation embeddings as a proxy for modelling rater’s annotation tendencies, as seen in Deng et al. [37]. In the case of AART [31], InfoNCE loss is incorporated to bring annotators with similar annotation patterns closer together, thereby improving model performance. Although these approaches have demonstrated better performance compared to single-task models trained on aggregated labels, a key limitation we have observed across all methods modelling individual perspectives is the restricted number of annotations delivered per annotator. Additionally, scalability becomes a challenge when dealing with hundreds or even thousands of annotators.
Secondly, modelling disagreement among groups of annotators is a learning approach that, instead of focusing solely on individual annotators, groups them based on shared characteristics such as religion, gender, political views, or ethnicity, among others. As observed in Fleisig et al. [38], this methodology relies on the idea that individuals with similar backgrounds may share annotation biases that can provide valuable contextual information and improve the predictive performance of individual annotations. One method for incorporating this information is by introducing the characteristics of the annotator as a prior textual prompt to the model’s prediction process using the token separator [SEP], as was done in Fleisig et al. [38]. A similar approach is taken by Wan et al. [39], with an additional step: training the model on labels aggregated within similar annotator groups. A distinct scheme is proposed by Gordon et al. [40], which reframes text classification as a recommendation problem. Their model integrates text embeddings, annotator embeddings, and annotator characteristic embeddings within a cross network and deep network architecture. This structure enables the model to learn personalized classification patterns based on annotator tendencies. However, despite these advantages, these approaches face several challenges. Among them, one major limitation is the difficulty in obtaining personal data characterizing the annotators, as such information is often sensitive or even unavailable. Additionally, individuals who share general demographic traits may still exhibit diverse annotation biases, as personal experiences significantly influence the decisions during the annotation process.
The third category, modelling label distribution, involves capturing the distribution of labels in a disaggregated dataset without explicitly modelling individual or group annotator perspectives. Several methods have been proposed to model label distribution. For instance, Plank et al. [41] introduced an inter-annotation agreement loss in the Part-of-Speech (POS) tagging task, leading to improved performance in downstream tasks. Similarly, Xu et al. [42] reformulated the task as a multilabel classification problem to model the full label distribution provided by annotators. Other methods consider label distribution as an auxiliary task. For example, Fornaciari et al. [11] predict soft labels, while Sandri et al. [29] convert the auxiliary task into classifying disagreement levels as none, low, or high. In Rodr´ıguez-Barroso et al. [43], the authors propose using Federated Learning to jointly train a model across different annotators without explicitly relying on majority voting. This approach enhances robustness and represents more fairly the plurality of opinions. Similarly to the approach proposed in this manuscript, Pavlopoulos and Likas [44] treat disagreement as a separate class to be predicted, incorporating the Distance from Unimodality (DFU) metric, which improves predictive performance for other labels. These approaches demonstrate the benefits of leveraging annotation disagreement to enhance the primary task’s performance. However, they face challenges due to the limited number of annotators per instance, making them susceptible to noise from ambiguous, inconsistent, or incomplete annotations.
@@ 2.3 Content Moderation
Before the rise of social media, CM was primarily performed by human moderators. However, with the surge of interactions on new platforms, this approach became unfeasible for humans to manage alone. As a result, companies like YouTube, Twitter, and Facebook began to develop their own moderation tools. Other initiatives also emerged, such as Google’s Perspective API and OpenAI’s CM tools, to support companies that did not have the resources to create their own tools. These algorithmic moderation tools were primarily designed to automate CM, often without fully considering the role of human moderators. However, this system has been observed to have several limitations regarding biases, including issues with state-of-the-art toxicity detectors like the Perspective API, as well as very limited capabilities in detecting deliberately obfuscated toxic sentences [46].
@@ 2.4 Main Contribution
Taking into account the analysed literature, the main contribution of this manuscript is a collaborative CM framework where the moderator is in control of the algorithmic CM tools by deciding when the moderator’s final decision should be incorporated. This is achieved not only by considering the model’s confidence through UQ, but also by incorporating information about the level of annotation disagreement the model should handle in its predictions. Since toxicity is a subjective task, with opinions varying across content annotators, it is crucial to leave the final judgment to those who are familiar with CM policies in place, as they are the ones who best understand these policies and their context of application. Additionally, given the evolving nature of language, it is of utmost importance to have mechanisms that are aware of this variability, especially in what refers to the ambiguity of textual toxicity. Along this line, the collaborative CM framework proposed in this work introduces information about the annotation disagreement into the modelling process within a multitask learning setting, enriching the conformalized estimation of the model’s confidence, and ultimately yielding better and more reliable decisions about the toxicity of the text data fed at its input. Our approach differs from existing annotation disagreement multitask learning methods used for modeling label distribution not only by leveraging annotation disagreement to improve the primary task (namely, toxicity classification), but also by transforming the auxiliary task into an interpretable variable that can directly inform correspondingly, the need for revising the comment flagged as toxic. In addition, we explore the incorporation of annotation disagreement within the uncertainty estimation paradigm, enhancing not only the performance but also the confidence calibration of the model. As demonstrated in Section 5, our multitask architecture achieves competitive detection performance, better-aligned and well-calibrated confidence estimations, and additional predictive insights (such as annotation ambiguity), enabling more informed moderation decisions for potentially toxic content.
@@ 3.1 Dataset
The Jigsaw Unintended Bias in Toxicity Classification dataset, created as part of a challenge organized by Google Jigsaw, provides a high-quality resource that is pre-divided into training, validation, and test subsets. This predefined partitioning enhances the reproducibility of experiments held with this dataset. The dataset contains comments from the archive of the Civil Comments platform, spanning from 2015 to 2017. Compared to other datasets released in the toxicity research area [4], this dataset includes a significantly larger group of annotators (8,899) in contrast to the smaller groups of tens typically found in other repositories. This large number of annotators offers a substantial advantage, as it provides a more diverse range of opinions across different backgrounds, ethnicities, genders, and ages. Furthermore, each comment is annotated by at least three annotators, with some comments receiving annotations from thousands of annotators. This large number of annotations helps reduce the likelihood of noisy signals resulting from inconsistent labelling, vagueness, or incomplete information and provides a more adequate representation of disagreement, as highlighted in Cabitza et al. [34] and Aroyo and Welty [33]. However, during the dataset preprocessing phase, several caveats were identified:
@@ 3.2 Methods
In this section we introduce the notation, selected neural network architecture, learning objectives, and hyperparameters (Subsection 3.2.1), whereas Subsection 3.2.2 discusses various methods from the literature for computing disagreement and explain our selected approach. Finally, Subsection 3.2.3 presents the CP algorithms considered in our experiments.
3.2.1 Notation, Neural Network Architecture, Learning Objectives and Hyperparameters
@@
As depicted in Figure 1, the MTL CM process consists of a primary task (toxicity classification of the given comment $\mathbf{x}_{i}$) and an auxiliary task (estimation of the annotation disagreement). The neural architecture is based on $\text{DistillBERT}_{\text{base uncased}}$ [51], which serves to extract the embedding from the CLS token (a sentence-level representation for classification) of the given comment $\mathbf{x}_{i}$ to perform the MTL task. The generated predictions for each task are defined as $\widehat{y}_{i}$ from the primary task and $\widehat{d}_{i}(\mathbf{x}_{i})$ from the auxiliary task. We then generate the corresponding prediction set $\mathcal{C}(\mathbf{x}_{i})$ and confidence interval $I(\mathbf{x}_{i})$ using the inductive CP techniques later detailed in Subsection 2.1 and a calibration subset $\mathcal{D}_{\text{cal}}$. Finally, the need for a human moderator’s revision is determined by whether prediction set length of $\mathcal{C}(\mathbf{x}_{i})$ is 2 (namely, the number of possible classes in our problem) or if the predicted confidence interval $I(\mathbf{x}_{i})$ is larger than an ambiguity threshold $\gamma$. The ambiguity threshold $\gamma$ quantifies the level of disagreement assigned to a comment during the annotation process, where lower values indicate stronger consensus among annotators, and higher values reflect greater divergence in opinions. This criterion allows considering the model’s inherent uncertainty in the predictions of the primary and auxiliary tasks, as well as the ambiguity of the comments. By incorporating this threshold, the framework enhances moderation efficiency by reducing the cognitive burden on moderators, allowing them to allocate more attention to complex or novel instances of toxic content.
@@ 3.2.3 Conformal Algorithms under Consideration
  • Class-Conditional LAC (CCLAC) [24]: In this case, we compute the conformity score $s(\mathbf{x},y)$ for each sample $\mathbf{x}$ and class $y$ as one minus the softmax probability assigned by the model to the class, either toxic or non-toxic. The quantile is then computed for each class, yielding a quantile value $\widehat{q}_{y}$ for each class $y$ (toxic and not toxic). Finally, during inference we iterate through the classes and include them in the prediction set associated to query instance $\mathbf{x}_{i}$ if their conformity score $s(\mathbf{x}_{i},y)$ is below the corresponding quantile $\widehat{q}_{y}$, namely:
@@
[…]
@@
[…]